If M is a singular Matrix, is there a value of k∈N for which kM will be non singular? Give reasons for your answer
1
Expert's answer
2020-12-08T05:12:52-0500

A square matrix is called singular if its determinant is 0. So, "\det (M)=0". Taking into account that "\det (kM)=k^n\det (M)" where "n" is a number of rows in "M", we conclude that "\det (kM)=k^n\det (M)=k^n\cdot 0=0" for all "k\in \mathbb N." Therefore, "kM" is a singular matrix as well.